You make a deposit of $927 into a savings account at the end of each year for the next 9 years. How much can you withdraw immediately after your last deposit if your savings account pays 7% compounded annually? Enter your answer as follows: 12345 Round your answer. Do not use a dollar sign ($), commas() or a decimal point (
Answer #1

Solution:

Amount = 927
9 payments will be made
1 at the end of each year
Intrest rate = 7% = 0.07
Formula for compound intrest:
A=P(1+r)^t
Amount to be recieved after 9th payment=
927(1.07)8 + 927(1.07)7 + 927(1.07)6 + 927(1.07)5 + 927(1.07)4 + 927(1.07)3 + 927(1.07)2 + 927(1.07) + 927(1.07)0
=1592.75 + 1488.55 + 1391.17 + 1300.16 +1215.10 +1135.61 +1061.32 + 991.89 + 927== 11103
Amount that can be withdrawn immediately after last deposit
11103 ans.